On a test given to all high school students in the state of Maryland, the average score was 82 and the standard deviation was 4.2. If a sample of 100 students is randomly selected, what is the probability that the sample's average is less than 83?
Answer by stanbon(75887)

On a test given to all high school students in the state of Maryland, the average score was 82 and the standard deviation was 4.2. If a sample of 100 students is randomly selected, what is the probability that the sample's average is less than 83?
---
Determine the z-score of 83.
---
z(83) = (83-82)/[4.2/sqrt(100)] = 1/[0.42] = 2.3810
----
P(xbar < 83) = P(z < 2.3810) = 0.9914
======================
Cheers,
Stan H.
======================